English

Solve the following: Find the probability that a year selected will have 53 Wednesdays

Advertisements
Advertisements

Question

Solve the following:

Find the probability that a year selected will have 53 Wednesdays

Sum
Advertisements

Solution

Let A ≡ the event that a year selected has 53 Wednesdays

L ≡ the event that leap year is selected

N ≡ the event that non-leap year is selected

The required event will happen if any one of L ∩ A and N ∩ A occurs.

These events are mutually exclusive

∴ the required probability

= P(L ∩ A) + P(N ∩ A)

= `"P"("L")*"P"("A"/"L") + "P"("N")*"P"("A"/"N")`  ...(1)

There is one leap year in 4 consecutive years

∴ P(L) = `1/4`

`"P"("A"/"L")` = Probability that year has 53 Wednesdays given that it is a leap year

Leap year has 366 day

366 = 7 x 52 + 2

∴ there are 52 full weeks and 2 days.

These days can be Sunday, Monday; Monday, Tuesday; Tuesday, Wednesday; Wednesday, Thursday; Thursday, Friday; Friday, Saturday; Saturday, Sunday.

There are 7 possibilities and favourable cases are 2

∴ `"P"("A"/"L") = 2/7`

Since there are 3 non-leap years in 4 consecutive years,

P(N) = `3/4`

`"P"("A"/"N")` = Probability that year has 53 Wednesdays given that it is a non-leap year

Non-leap year has 365 days

365 = 7 x 52 + 1

∴ there are 52 full weeks and 1 day.

This day could be any day of the week days i.e., any one of 7 days.

The number of favourable case is 1.

∴ `"P"("A"/"N") = 1/7`

∴ from (1), the required probability = `1/4*2/7 + 3/4*1/7`

= `5/28`.

shaalaa.com
  Is there an error in this question or solution?
Chapter 9: Probability - Miscellaneous Exercise 9 [Page 214]

APPEARS IN

Video TutorialsVIEW ALL [2]

RELATED QUESTIONS

A card from a pack of 52 playing cards is lost. From the remaining cards of the pack three cards are drawn at random (without replacement) and are found to be all spades. Find the probability of the lost card being a spade.


If `P(A)  = 3/5 and P(B) = 1/5` , find P (A ∩ B) if A and B are independent events.


A fair coin and an unbiased die are tossed. Let A be the event ‘head appears on the coin’ and B be the event ‘3 on the die’. Check whether A and B are independent events or not.


If A and B are two events such that `P(A) = 1/4, P(B) = 1/2 and P(A ∩ B) = 1/8`, find P (not A and not B).


A problem in statistics is given to three students A, B, and C. Their chances of solving the problem are `1/3`, `1/4`, and `1/5` respectively. If all of them try independently, what is the probability that, problem is not solved


The odds against a certain event are 5: 2 and odds in favour of another independent event are 6: 5. Find the chance that at least one of the events will happen.


The odds against student X solving a business statistics problem are 8: 6 and odds in favour of student Y solving the same problem are 14: 16 What is the probability that neither solves the problem?


The probability that a student X solves a problem in dynamics is `2/5` and the probability that student Y solves the same problem is `1/4`. What is the probability that

  1. the problem is not solved
  2. the problem is solved
  3. the problem is solved exactly by one of them

Two hundred patients who had either Eye surgery or Throat surgery were asked whether they were satisfied or unsatisfied regarding the result of their surgery.

The following table summarizes their response:

Surgery Satisfied Unsatisfied Total
Throat 70 25 95
Eye 90 15 105
Total 160 40 200

If one person from the 200 patients is selected at random, determine the probability the person had Throat surgery given that the person was unsatisfied.


The probability that a man who is 45 years old will be alive till he becomes 70 is `5/12`. The probability that his wife who is 40 years old will be alive till she becomes 65 is `3/8`. What is the probability that, 25 years hence,

  1. the couple will be alive
  2. exactly one of them will be alive
  3. none of them will be alive
  4. at least one of them will be alive

A family has two children. Find the probability that both the children are girls, given that atleast one of them is a girl.


Solve the following:

If P(A) = `"P"("A"/"B") = 1/5, "P"("B"/"A") = 1/3` the find `"P"("B'"/"A'")`


Solve the following:

Let A and B be independent events with P(A) = `1/4`, and P(A ∪ B) = 2P(B) – P(A). Find `"P"("B'"/"A")`


Solve the following:

For three events A, B and C, we know that A and C are independent, B and C are independent, A and B are disjoint, P(A ∪ C) = `2/3`, P(B ∪ C) = `3/4`, P(A ∪ B ∪ C) = `11/12`. Find P(A), P(B) and P(C)


Solve the following:

A and B throw a die alternatively till one of them gets a 3 and wins the game. Find the respective probabilities of winning. (Assuming A begins the game)


10% of the bulbs produced in a factory are of red colour and 2% are red and defective. If one bulb is picked up at random, determine the probability of its being defective if it is red.


Two dice are thrown together. Let A be the event ‘getting 6 on the first die’ and B be the event ‘getting 2 on the second die’. Are the events A and B independent?


If A and B are independent events such that P(A) = p, P(B) = 2p and P(Exactly one of A, B) = `5/9`, then p = ______.


For a loaded die, the probabilities of outcomes are given as under:
P(1) = P(2) = 0.2, P(3) = P(5) = P(6) = 0.1 and P(4) = 0.3. The die is thrown two times. Let A and B be the events, ‘same number each time’, and ‘a total score is 10 or more’, respectively. Determine whether or not A and B are independent.


A and B are two events such that P(A) = `1/2`, P(B) = `1/3` and P(A ∩ B) = `1/4`. Find: `"P"("A"/"B")`


A and B are two events such that P(A) = `1/2`, P(B) = `1/3` and P(A ∩ B) = `1/4`. Find: `"P"("B"/"A")`


Let E1 and E2 be two independent events such that P(E1) = P1 and P(E2) = P2. Describe in words of the events whose probabilities are: 1 – (1 – P1)(1 – P2) 


Two dice are tossed. Find whether the following two events A and B are independent: A = {(x, y): x + y = 11} B = {(x, y): x ≠ 5} where (x, y) denotes a typical sample point.


If A and B are two events and A ≠ Φ, B ≠ Φ, then ______.


If two events are independent, then ______.


Two events E and F are independent. If P(E) = 0.3, P(E ∪ F) = 0.5, then P(E|F) – P(F|E) equals ______.


If A and B are mutually exclusive events, then they will be independent also.


If A and B′ are independent events, then P(A' ∪ B) = 1 – P (A) P(B')


Two events 'A' and 'B' are said to be independent if


If P(A) = `3/5` and P(B) = `1/5`, find P(A ∩ B), If A and B are independent events.


Given two independent events A and B such that P(A) = 0.3, P(B) = 0.6 and P(A' ∩ B') is ______.


Let EC denote the complement of an event E. Let E1, E2 and E3 be any pairwise independent events with P(E1) > 0 and P(E1 ∩ E2 ∩ E3) = 0. Then `"P"(("E"_2^"C"  ∩ "E"_3^"C")/"E"_1)` is equal to ______.


A die is thrown once. What is the sample space \[S\]?


A die is thrown once. If E is the event that the number appearing is a multiple of 3, which set represents E?


For \[E=\{3,6\}\] and \[F=\{2,4,6\}\], what is \[E\cap F\]?


For \[E=\{3,6\}\] when a die is thrown once, what is \[P(E)\]?


For \[F=\{2,4,6\}\] when a die is thrown once, what is \[P(F)\]?


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×